Suppose Honda’s total cost of producing 4 cars is $225,000 and its total cost of producing 5 cars is $250,000. What is the average total cost of producing 5 cars? What is the marginal cost of the fifth car?  Draw the marginal-cost curve and the average-total-cost curve for a typical firm, and explain why these curves cross where they do.



1
Expert's answer
2021-06-07T06:12:06-0400

The average cost of producing 5 cars

"Average" total cost "=(\frac{Total\space cost}{Quantity})"


"=(\frac{250,000}{5})=50,000"


Thus the value of average total cost is $50,000


The marginal cost of the fifth car can be calculated as


Marginal cost "=(\frac{Total \space cost_5\space - Total \space cost_4}{5-4})"


"=(\frac{250,000-225,000}{5-4})"


"=25,000"


Thus the marginal cos for the fifth car is $25,000


Marginal Cost curve of a typical firm




In the figure above, the horizontal axis measures the quantity and the vertical axis represents the cost. The curves cross at the efficient scale because at the low level of output, the marginal cost curve is below the average total cost and it leads to a fall in average total cost. After the intersection of both the curves, marginal cost starts increasing above the average total cost. Thus, the intersection point must be at minimum of average total cost